OpenStudy (anonymous):

The solution to the radical function is x = 8.Answer True False

OpenStudy (anonymous):



OpenStudy (anonymous):

@jennychan12

OpenStudy (anonymous):

\[\sqrt{x+3} = 2\] \[x+3 = 2^{2}\] x+3= 4 x= 4-3 x =1

OpenStudy (jennychan12):

square both sides x+3 = 4 x = 1

OpenStudy (jennychan12):

false.

OpenStudy (anonymous):

True or false? so false